Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 33     Entry time: 15 Feb 2023, 23:05
Author: Luigi, Laurent 
Group: SCS 
Subject: late shift summary 2nd day 

Achievememts:

  1. optimised the beam rep rate to value of 140KHz for a better signal 
  2. changed nozzles and refilled the reservoir
  3. Improved notebook for data analysis
  4. Cammille made an online tool to see correlations between Viking and XGM --> very useful!

Night shift:

  1. repeating measurements with one color with the new settings
  2. following for the 2 colors plan

 

ELOG V3.1.4-7c3fd00